I put 1.422*10^-5 for part A--incorrec

for part b i put 7.931*10^-5

for c i put 3.22*10^-4

all were incorrect, please help

Item 9 A single-turn square loop of side Lis cente
Answer #1


N =1160 turns, d = 6.45 cm , r = 3.225 cm,I = 2.65 A

B = uoNI/2pi*r

flux phi = BA = (uoNI/2pi*r)*L^2

(A) L = 2.75 cm

phi = (2*10^-7*1160*2.65/(2*3.14*0.03225))*0.0275^2

phi = 2.3*10^-6 Wb

(B) L = 6.45 cm

phi = (2*10^-7*1160*2.65/(2*3.14*0.03225))*0.0645^2

phi = 1.263*10^-5 Wb

(C)L =13 cm

phi = (2*10^-7*1160*2.65/(2*3.14*0.03225))*0.13^2

phi = 5.13*10^-5 Wb